OpenStudy (anonymous):

Solve (x + 2 < 5) U (x - 7 > -6). {x | -6 < x < 5} {x | x < 3 or x > 1} Ø

OpenStudy (anonymous):

a^2+2ab+c^2 .... so now here 2ab=2*5*1 and now link a which is 1 so be is 5 so add some to make 7 to 25 add 18 both sides then (x+5)^2 you will get (x+5)^2=18

OpenStudy (anonymous):

U means or x+2<5 or x-7>-6 x<3 or x>1 that is for all x in real numbers, the statement is true

OpenStudy (anonymous):

The correct procedure for "completing the square" is as follows: fir shift the constant to the other side x^2 + 10x = -7 Now add the square of half of the coefficient of x to both sides here coefficient of x is 10, its half is 5, so add 5^2 to both sides x^2 + 10x + 5^2 = -7 + 5^2 Now LHS become (x+5)^2, so we can write (x+5)^2 = -7 + 25 taking root of both sides x+5 = +/-root(18) x = -5 +/- 3root2 x = -5 + 3root2 or x = -5 - 3root2

OpenStudy (anonymous):

completing the square not finding the value of x ??

OpenStudy (anonymous):

the question said "solve by completing the square" so i guess we have to find values of x....

OpenStudy (anonymous):

so B

OpenStudy (anonymous):

@jonjenkins7653

OpenStudy (anonymous):

yes